OpenStudy (anonymous):

PLLLEEEZZZ HELPP!!! FAST! Simplify completely ((x2-3x-54)/(x2-18x+81)) * ((x2+12x+36)/(x+6))

OpenStudy (anonymous):

factor x^2 - 3x - 54 factor x^2 - 18x + 81 factor x^2 + 12x + 36 ~

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

x^2 - 3x - 54 = x^2 - 9x + 6x - 54 = (x-9) (x+6) x^2 - 18x + 81 = (x-9)^2 x^2 + 12x + 36 = (x+6)^2 you know how to go on? o.o

OpenStudy (anonymous):

yah/..... whats the answer?

OpenStudy (anonymous):

\[\frac{(x-9)(x+6)}{(x-9)^2}\times \frac{(x+6)^2}{(x+6)}\]\[\frac{(x+6)}{(x-9)}\times (x+6)\]\[\frac{(x+6)^2}{(x-9)}\]